# All or None: Identifiable Linear Properties of Next-token Predictors in Language Modeling

### v1

@@ 4.2.3 Linear Steering
Similar to LP, in Section C.1 we discuss a relational counterpart to linear properties discussed by Park et al. [23], Hernandez et al. [2].
@@ 5 Linear properties shared by all distribution-equivalent models
Based on Theorem 5, we analyze which linear properties are shared across models expressing the same conditional distribution. To this end, pick a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ that linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma$ (Definition 7), and also consider the space $\Gamma_{\bm{\mathrm{q}}}:=\mathrm{Im}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})$. We show that under an additional condition on $\Gamma$ and $\Gamma_{\bm{\mathrm{q}}}$, two models that are $\sim_{EL}$-equivalent share the same linear properties (all proofs are in Appendix D):
Proposition 14.
@@
For two models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}),(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})\in\Theta$, such that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$, it holds: if ${\bm{\mathrm{f}}}$ linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma\subseteq\mathcal{N}$, and moreover, $\Gamma_{\bm{\mathrm{q}}}\subseteq\mathcal{M}$, then $\tilde{\bm{\mathrm{f}}}$ linearly represents ${\bm{\mathrm{q}}}$ on $\tilde{\Gamma}\subseteq\tilde{}\mathcal{N}$, where $\tilde{\Gamma}=\mathrm{Im}({\bm{\mathrm{N}}}^{+}{\bm{\mathrm{P}}}_{\Gamma})$ and ${\bm{\mathrm{N}}}$ is the matrix relating ${\bm{\mathrm{g}}}_{0}$ and $\tilde{{\bm{\mathrm{g}}}}_{0}$ by the equivalence relation in Definition 3.
@@
This shows how the relational linearity ($\Gamma$lr), combined with the condition that $\Gamma\subseteq\mathcal{N}$ and $\Gamma_{\bm{\mathrm{q}}}\subseteq\mathcal{M}$ is a property of all or none next-token predictors modeling the same conditional distribution. As a consequence, the same holds for ls (by Proposition 11) and lp (by Proposition 13). Intuitively, the extra condition underlies that relational linearity of $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ is displayed by the components of ${\bm{\mathrm{f}}}$ that contribute to the dot product with ${\bm{\mathrm{g}}}_{0}$. A $\sim_{EL}$-equivalent model $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ would linearly transform these components, thus preserving relational linearity. Vice versa, since all components of ${\bm{\mathrm{f}}}$ outside $\mathcal{M}$ can be arbitrarily distorted, any property of $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ that depends on those components may not hold for $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$. The extra condition precisely avoids that. [^3] If $\Gamma\not\subseteq\mathcal{N}$, then relational linearity (Definition 7) would be trivially satisfied for $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ and, in turn, also for $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$, because ${\bm{\mathrm{\gamma}}}\in\Gamma$ would give ${\bm{\mathrm{\gamma}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})=0$, and so ${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})=\mathbf{0}$. Notice that the special case where the diversity condition (Definition 1) holds implies a similar conclusion because the condition that $\Gamma\subseteq\mathcal{N}$ and $\Gamma_{\bm{\mathrm{q}}}\subseteq\mathcal{M}$ is then always satisfied (as $\mathcal{M}=\mathcal{N}=\mathbb{R}^{d}$).
In contrast, parallel vectors may not be preserved: Two parallel vectors in one model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ may not be parallel in another model $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ with the same conditional distribution. They remain parallel only within the subspaces $\mathcal{N}$ and ~$\tilde{}\mathcal{N}$, respectively:
Proposition 15.
@@ 6 Discussion
Proposition 14 gives an example of a property that all distribution-equivalent next-token predictors, as characterized by our identifiability result (Theorem 5), must share. For the widely observed linear properties of language models, one may then ask whether they are indeed examples of shared properties akin to the one in Proposition 14. Tautologically, claims about the ubiquity of linear properties cannot solely be based on observations of linearity in individual model instances. One might therefore expect that only those properties shared across all equivalent models should be ubiquitously observable. [^4] An analogy could be made with the principle of covariance in physics [32, 33], which asserts that physical laws should be expressible as coordinate-independent and reference-frame-independent geometric relationships between objects that represent physical entities [34]. Recently, Villar et al. [35] suggested that this principle could inspire future developments in machine learning. In the following, we critically examine this hypothesis in light of empirical evidence.
@@
Proposition 15 shows that vector parallelism is only preserved within a linear subspace $\mathcal{N}$ of the unembedding space. Surprisingly, a different kind of parallelism, which according to our theory is not shared by all $\sim_{EL}$-equivalent models, appears to be consistently observed in language models. In fact, several empirical studies apply dimensionality reduction through PCA to the embeddings and unembeddings to reveal and visualize linear properties including parallelism (e.g., [9, Figure 2]; [15, Figure 1]). Note that, across $\sim_{EL}$-equivalent models, the embeddings and unembeddings may not be completely contained within $\mathcal{M}$ and $\mathcal{N}$, respectively: that is the case when both $\mathcal{M}\subsetneq\mathrm{SIm}({\bm{\mathrm{f}}})$ and $\mathcal{N}\subsetneq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$. As a consequence, unbounded distortions within the orthogonal complements of $\mathcal{M}$ and of $\mathcal{N}$ are inconsequential for the dot product between the embedding and the unembedding vectors (see, e.g.Figure 2, top left, for an embedding manifold not contained within any proper linear subspace). If these distortions were sufficiently large, they would prevent the visualization of vectors parallel in the sense of Proposition 15 through PCA, as the distortions would dominate the covariance matrix on which the PCA of the representations is performed, and thus the first principal components would mostly reflect those. This suggests that, in models where PCA reveals parallelism, these distortions are small, and the representations live close to a proper linear subspace.
@@ 7 Related work and future directions
Linear properties of next-token predictors have attracted widespread attention, also beyond language modeling [36, 37, 38]. More complex, non-linear properties have also been observed, such as circular token representations [39]. Formalizing these properties and investigating whether all distribution-equivalent next-token predictors share them, in the sense we studied for linear properties, is an interesting open venue.
Theoretical studies on linear properties. Park et al. [23] introduce binary latent concepts to describe several linear properties (though not relational linearity) in a unified framework. This was also applied to study categorical and hierarchical concepts [40]. Jiang et al. [28] explain linear properties of language models based on assumptions on the data-generating process and latent variables underlying natural text. This allows them to reason about the origins of linearity; in this work, we instead focus on the ubiquity of linearity in the models, taking an agnostic stance on latent concept variables. An exciting direction for future work is to prove, within our framework, why and how linear properties emerge, if they do at all.
Identifiability of representations is a central theme in generative modeling, [41, 42], particularly in non-linear ICA, e.g., [43, 44, 45, 46, 47]; and in causal representation learning, e.g., [48, 49, 50, 51, 52, 53, 54]. Reizinger et al. [55] discuss the role identifiability may play in explaining several properties of large language models [56]. Our work highlights the role identifiability plays in explaining the ubiquity of linear properties.
@@ Acknowledgments
We thank Beatrix Miranda Nielsen, Antonio Vergari, Frederik Hytting Jørgensen, Filippo Camilloni, Adrián Javaloy, and Julius von Kügelgen for useful discussions. We acknowledge positive feedback by Stefano Teso and interesting conversations with the participants of the 2024 Bellairs Workshop on Causality. E.M. acknowledges support from TANGO, Grant Agreement No. 101120763. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. L.G. was supported by the Danish Data Science Academy (DDSA), Grant ID: 2023-1250.
@@ Appendix A Functional Form of Decoders-only Transformers
Decoders-only transformer models can be reduced to the form we use in the main text, as already shown in a derivation by Roeder et al. [3]. We consider autoregressive, GPT-like models [57, 58], focusing on the GPT-J model [59].
@@
This explicit form of the embedding and unembedding respectively and consistent with (1), as previously detailed by Roeder et al. [3].
@@ B.1 Reminder of useful properties of pseudo-inverses
We will often make use of the pseudo-inverse ${\bm{\mathrm{A}}}^{+}$ of a matrix ${\bm{\mathrm{A}}}$ [60, page 221]. We denote with ${\bm{\mathrm{T}}}\mid_{\mathrm{ker}({\bm{\mathrm{T}}})^{\perp}}$ the restriction of ${\bm{\mathrm{T}}}$ to its orthogonal complement of the kernel [60].
@@ B.2 Proof of Lemma 2
(i) By the rank-nullity theorem [60, page 62], we have that $\mathrm{dim}\,\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})=d-\mathrm{dim}\,\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})$. Notice that $\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})=\mathcal{M}$ and $\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})=\mathbb{R}^{d}\setminus\mathcal{N}$, therefore
@@
Next, we derive the dimensionality of $\mathcal{M}$. Recall that for two matrices ${\bm{\mathrm{A}}}\in\mathbb{R}^{m\times k}$ and ${\bm{\mathrm{B}}}\in\mathbb{R}^{k\times n}$, from the rank-nullity theorem [60, page 62], it follows that:
@@
(ii) From the property of the pseudo-inverse, see 27 and 28, we have that
@@
From 45, taking the transpose of ${\bm{\mathrm{P}}}_{\mathcal{M}}$ we get:
@@
and from 46, taking the transpose we obtain:
@@
(iii) Denote ${\bm{\mathrm{A}}}:={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}$. From the pseudo-inverse definition [60, page 221], it holds
@@
Similarly to point (ii), taking the transpose of 54, we obtain:
@@
and taking the transpose of 56 we obtain:

### v2

@@ 4.2.3 Linear Steering
Another property that has attracted considerable attention is the linear steering property [Stolfo et al., 2024], also termed linear intervening property by Park et al. [2024a]. By knowing what queries are linearly represented by the model (as per Definition 9), this property allows us to steer the model embeddings such that the most-likely reply to a given query changes, while the replies to other queries remain unaffected. In Section C.1, we define a relational version of this property, and show under what conditions it is implied by the relational linear property (Proposition 19).
@@ 5 Linear properties shared by all distribution-equivalent models
Based on Theorem 5, we can now analyze which linear properties are shared across models expressing the same next-token distribution. We start from relational linearity (Definition 9). To this end, pick a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ that linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma$, and consider the space $\Gamma_{\bm{\mathrm{q}}}:=\mathrm{Im}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})$. We show that under an additional condition on $\Gamma$ and $\Gamma_{\bm{\mathrm{q}}}$, two models that are $\sim_{EL}$-equivalent share the same linear properties (results from this section are proved in Appendix D):
Theorem 14.
@@
For two models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}),(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})\in\Theta$ s.t. $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$, if ${\bm{\mathrm{f}}}$ linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma\subseteq\mathcal{N}$, and $\Gamma_{\bm{\mathrm{q}}}\subseteq\mathcal{M}$, then $\tilde{\bm{\mathrm{f}}}$ linearly represents ${\bm{\mathrm{q}}}$ on $\tilde{\Gamma}\subseteq\tilde{}\mathcal{N}$, where $\tilde{\Gamma}=\mathrm{Im}({\bm{\mathrm{N}}}^{+}{\bm{\mathrm{P}}}_{\Gamma})$ and ${\bm{\mathrm{N}}}$ is the matrix relating ${\bm{\mathrm{g}}}_{0}$ and $\tilde{{\bm{\mathrm{g}}}}_{0}$ by the equivalence relation in Definition 3.
@@
This shows how, under the the condition that $\Gamma\subseteq\mathcal{N}$ and $\Gamma_{\bm{\mathrm{q}}}\subseteq\mathcal{M}$, relational linearity ($\Gamma$lr) is a property of all or none next-token predictors modeling the same conditional distribution. As a consequence, the same holds for ls (by Proposition 11) and lp (by Proposition 13). Intuitively, the extra condition underlies that relational linearity of $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ is displayed by the components of ${\bm{\mathrm{f}}}$ that contribute to the dot product with ${\bm{\mathrm{g}}}_{0}$. A $\sim_{EL}$-equivalent model $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ would linearly transform these components, thus preserving relational linearity. Vice versa, since all components of ${\bm{\mathrm{f}}}$ outside $\mathcal{M}$ can be arbitrarily distorted, any property of $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ that depends on those components may not hold for $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$. The extra condition precisely avoids that. [^5] If $\Gamma\not\subseteq\mathcal{N}$, then relational linearity (Definition 9) would be trivially satisfied for $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ and, in turn, also for $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$, because ${\bm{\mathrm{\gamma}}}\in\Gamma$ would give ${\bm{\mathrm{\gamma}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})=0$, and so ${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})=\mathbf{0}$. Notice that the special case where the diversity condition (Definition 1) holds implies a similar conclusion because the condition that $\Gamma\subseteq\mathcal{N}$ and $\Gamma_{\bm{\mathrm{q}}}\subseteq\mathcal{M}$ is then always satisfied (as $\mathcal{M}=\mathcal{N}=\mathbb{R}^{d}$). This testifies that (a special case of) relational linearity is shared among $\sim_{EL}$ models. In contrast, vector parallelism may not be preserved: Two parallel vectors in one model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ may not be parallel in another model $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ with the same conditional distribution. They remain parallel only within the subspaces $\mathcal{N}$ and ~$\tilde{}\mathcal{N}$, respectively:
Theorem 15.
@@ 6 Discussion
Theorem 14 is an example of a property that all distribution-equivalent next-token predictors, as characterized by our identifiability result (Theorem 5), must share. One may then ask whether the widely observed linear properties of language models are indeed examples of shared properties akin to the one in Theorem 14. Tautologically, claims about the ubiquity of linear properties cannot solely be based on observations of linearity in individual model instances. One might thus hypothesize that only those properties shared across all equivalent models should be ubiquitously observable. [^6] An analogy could be made with the principle of covariance in physics [Einstein, 1920, Thorne et al., 2000], which asserts that physical laws should be expressible as coordinate-independent and reference-frame-independent geometric relationships between objects that represent physical entities [Thorne and Blandford, 2017]. Recently, Villar et al. [2023] suggested that this principle could inspire future developments in machine learning. In the following, we critically examine this hypothesis in light of empirical evidence.
@@
Theorem 15 shows that vector parallelism is only preserved within a linear subspace $\mathcal{N}$ of the unembedding space. Surprisingly, a different kind of parallelism, which according to our theory is not shared by all $\sim_{EL}$-equivalent models, appears to be consistently observed in language models. In fact, several empirical studies apply dimensionality reduction through PCA to the embeddings and unembeddings to reveal and visualize linear properties including parallelism (e.g., [Mikolov et al., 2013a, Figure 2]; [Marks and Tegmark, 2023, Figure 1]). Note that, across $\sim_{EL}$-equivalent models, the embeddings and unembeddings may not be completely contained within $\mathcal{M}$ and $\mathcal{N}$, respectively: that is the case when both $\mathcal{M}\subsetneq\mathrm{SIm}({\bm{\mathrm{f}}})$ and $\mathcal{N}\subsetneq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$. As a consequence, unbounded distortions within the orthogonal complements of $\mathcal{M}$ and of $\mathcal{N}$ are inconsequential for the dot product between the embedding and the unembedding vectors (see, e.g.Figure 2, top left, for an embedding manifold not contained within any proper linear subspace). If these distortions were sufficiently large, they would prevent the visualization of vectors parallel in the sense of Theorem 15 through PCA, as the distortions would dominate the covariance matrix on which the PCA of the representations is performed, and thus the first principal components would mostly reflect those. This suggests that, in models where PCA reveals parallelism, these distortions are small, and the representations live close to a proper linear subspace.
@@ 7 Related work and future directions
Linear properties of next-token predictors have attracted widespread attention, also beyond language modeling [Li et al., 2022, Nanda et al., 2023, Elhage et al., 2022]. More complex, non-linear properties have also been observed, such as circular token representations [Engels et al., 2024]. Formalizing these properties and investigating whether all distribution-equivalent next-token predictors share them, in the sense we studied for linear properties, is an interesting open venue.
Theoretical studies on linear properties. Park et al. [2024a] introduce binary latent concepts to describe several linear properties (though not relational linearity) in a unified framework. This was also applied to study categorical and hierarchical concepts [Park et al., 2024b]. Jiang et al. [2024] explain linear properties of language models based on assumptions on the data-generating process and latent variables underlying natural text. This allows them to reason about the origins of linearity; in this work, we instead focus on the ubiquity of linear properties, with an agnostic stance on latent concept variables.
An exciting direction for future work is to prove, within our framework, why and how linear properties emerge, if they do at all.
Identifiability of representations is a central theme in generative modeling [Moran et al., 2022, Xi and Bloem-Reddy, 2023], particularly in non-linear ICA [Hyvarinen et al., 2019, Gresele et al., 2020, Hälvä and Hyvarinen, 2020, Buchholz et al., 2022, Hyttinen et al., 2022] and causal representation learning [Lippe et al., 2022, Ahuja et al., 2023, Liang et al., 2024, von Kügelgen et al., 2024, Varici et al., 2024, Zhang et al., 2024a, Rajendran et al., 2024, Li et al., 2024, Bortolotti et al., 2025]. Buchholz [2024] studied when token partitions can be identified from their interactions; Reizinger et al. [2024] discussed what role identifiability may play in explaining several aspects of large language models [Zhang et al., 2024b]. Our work highlights the role of identifiability in explaining the ubiquity of linear properties in language models.
@@ Acknowledgments
We thank Beatrix Miranda Nielsen, Antonio Vergari, Frederik Hytting Jørgensen, Filippo Camilloni, Adrián Javaloy, and Julius von Kügelgen for useful discussions. We acknowledge positive feedback by Stefano Teso and interesting conversations with the participants of the 2024 Bellairs Workshop on Causality. E.M. acknowledges support from TANGO, Grant Agreement No. 101120763. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. L.G. was supported by Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429).
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes/No/Not Applicable]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes/No/Not Applicable]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes/No/Not Applicable]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes/No/Not Applicable]
    2. (b)
      Complete proofs of all theoretical results. [Yes/No/Not Applicable]
    3. (c)
      Clear explanations of any assumptions. [Yes/No/Not Applicable]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes/No/Not Applicable]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes/No/Not Applicable]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes/No/Not Applicable]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes/No/Not Applicable]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes/No/Not Applicable]
    2. (b)
      The license information of the assets, if applicable. [Yes/No/Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes/No/Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Yes/No/Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Yes/No/Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Yes/No/Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Yes/No/Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Yes/No/Not Applicable]
[…]
@@ Appendix A Functional Form of Decoders-only Transformers
Decoders-only transformer models can be reduced to the form we use in the main text, as already shown in a derivation by Roeder et al. [2021]. We consider autoregressive, GPT-like models [Liu et al., 2018, Radford et al., 2021], focusing on the GPT-J model [Wang and Komatsuzaki, 2021].
@@
This explicit form of the embedding and unembedding respectively and consistent with Equation (1), as previously detailed by Roeder et al. [2021].
@@ B.1 Reminder of useful properties of pseudo-inverses
We will often make use of the pseudo-inverse ${\bm{\mathrm{A}}}^{+}$ of a matrix ${\bm{\mathrm{A}}}$ [Axler, 2015, page 221]. We denote with ${\bm{\mathrm{T}}}\mid_{\mathrm{ker}({\bm{\mathrm{T}}})^{\perp}}$ the restriction of ${\bm{\mathrm{T}}}$ to its orthogonal complement of the kernel [Axler, 2015].
@@ B.2 Proof of Lemma 2
(i) By the rank-nullity theorem [Axler, 2015, page 62], we have that $\mathrm{dim}\,\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})=d-\mathrm{dim}\,\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})$. Notice that $\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})=\mathcal{M}$ and $\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})=\mathbb{R}^{d}\setminus\mathcal{N}$, therefore
@@
Next, we derive the dimensionality of $\mathcal{M}$. Recall that for two matrices ${\bm{\mathrm{A}}}\in\mathbb{R}^{m\times k}$ and ${\bm{\mathrm{B}}}\in\mathbb{R}^{k\times n}$, from the rank-nullity theorem [Axler, 2015, page 62], it follows that:
@@
(ii) From the property of the pseudo-inverse, see Equation (29) and Equation (30), we have that
@@
From Equation (47), taking the transpose of ${\bm{\mathrm{P}}}_{\mathcal{M}}$ we get:
@@
and from Equation (48), taking the transpose we obtain:
@@
(iii) Denote ${\bm{\mathrm{A}}}:={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}$. From the pseudo-inverse definition [Axler, 2015, page 221], it holds
@@
Similarly to point (ii), taking the transpose of Equation (56), we obtain:
@@
and taking the transpose of Equation (58) we obtain: